Lin Hisa plans to invest $1,400 in two stocks: Consolidated Industries (CI) and Amalgamated Manufacturing (AM). Suppose that after 5 years Lin's CI stock does not change in value, but her AM stock triples, making her stocks worth $3,300. Write an equation that expresses this fact.: Lin Hisa plans to invest $1,400 in two stocks: Consolidated Industries (CI) and Amalgamated Manufacturing (AM). Suppose that after 5 years Lin's CI stock does not change in value, but her AM stock triples, making her stocks worth $3,300. Write an equation that expresses this fact.
Answer by gonzo(568)

let x = amount of money invested in CI today
let y = amount of money invested in AM today.
-----
total invested in both equals $1400.
x + y = $1400
-----
5 years from now CI stock doesn't change in value.
5 years from now AM stock triples.
total value of both stocks in 5 years is $3300.
x + 3y = $3300
-----
you have two equations that you want to solve simultaneously.
x + y = 1400
x + 3y = 3300
-----
if you subtract the second equation from the firsts equation, the unknown variable of x will disappear.
you get:
(x - x) + (3y - y) = (3300 - 1400)
this becomes:
2y = 1900
y = 950
-----
substituting 950 for y in the first equation gets:
x + 950 = 1400
subtract 950 from both sides of the equation to get:
x = 1400 - 950
which becomes:
x = 450.
-----
you now have a value for x and y that should be able to satisfy both equations.
-----
first equation has already been satisfied since that is the equation where we solved for x.
-----
second equation is:
x + 3y = 3300
substituting 450 for x and 950 for y, we get:
450 + (3*950) = 3300
simplifying:
450 + 2850 = 3300
3300 = 3300
equation is satisfied.
-----
answer is:
x = amount invested in CI stock = $450
y = amount invested in AM stock = $950
-----